OpenStudy (anonymous):

Which equations are equivalent to the given equation? 49= -7k Choose all answers that are correct A. 49÷ (-7) = -7k ÷ (-7) B. 49∗ (-7) = -7k ÷ (-7) C. 49∗ | - 1 | = -7k ∗ (-7) | -- | | 7 | D. 49∗ | -1 | = -7k ∗ | -1 | | -- | | --- | | 7 | | 7 |

OpenStudy (anonymous):

I would say that the correct answer choice is D. A. 49÷ (-7) = -7k ÷ (-7) -7=1k B. 49∗ (-7) = -7k ÷ (-7) -343=1k C. 49∗ | - 1 | = -7k ∗ (-7) | -- | | 7 | (49/1)(-1/7) = -49/7 = -7 (-7k)(-7) = 49k -7=49k While, D. D. 49∗ | -1 | = -7k ∗ | -1 | | -- | | --- | | 7 | | 7 | (49/1)(-1/7) = -49/7= -7 (-7k/1)(-1/7) =1k -7=1k